An object with an initial velocity of 20m/s accelerates uniformly at 5m/s² in the direction of its motion for a diatance of 10m.What is the final velocity of the object


1
Expert's answer
2021-02-08T18:42:25-0500
"v_f^2=v_i^2+2ad,""v=\sqrt{v_i^2+2ad},""v=\sqrt{(20\ \dfrac{m}{s})^2+2\cdot5\ \dfrac{m}{s^2}\cdot10\ m}=22.36\ \dfrac{m}{s}."